Choose the Dark Forest for your selection Experiment

 
 
Percent Dark Moths
Percent Light Moths
Sooty Forest
65%
35%
  1. What was the gene frequency for the m allele before selection in this simulation? (Remember p2+ 2pq+ q2= 1)
  1. What would be the genotype frequencies before selection?

MM:

Mm:

mm:

  1. What was the frequency of the m allele after selection?
Answer #1

Before selection =underroot 35/ 100=Underroot 0.35 =0.59

thus acc to hardy equillibrium

M= 1-0.59=0.41

mm=0.41x0.41=0.1681

Mm=2x0.41x0.59=0.4838

MM= 0.59x0.59=0.3481

after selection =0.1681+1/2x0.4838/100

= 0.41